OpenStudy (anonymous):

If cosθ + cos^(2) θ = 1, then sin^(12) θ + 3 sin^(10) θ + 3 sin^(8) θ+ sin^(6) θ + 2 sin^(4) θ + 2 sin^(2) θ – 2 =

OpenStudy (anonymous):

Your clue is: \[\cos{\theta} + \cos^2{\theta}= 1\] \[\cos{\theta}=\sin^2{\theta}\]

OpenStudy (anonymous):

but i have solved it my answer is not cmg

OpenStudy (anonymous):

answer optios are: (A) 0 (B) 1 (C) 2 (D) 3

OpenStudy (anonymous):

Show me your working.

OpenStudy (anonymous):

plse see the attachment and when i put the value , i got -5 / 64 answer

OpenStudy (anonymous):

You mixed up something in the middle.

OpenStudy (anonymous):

\[\sin^2{\theta}=\sqrt{1-\sin^2{\theta}}\]

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

but at the end , i m not getting it

OpenStudy (anonymous):

Is that equation equal to 0?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

i have to calculate the value of sin^(12) θ + 3 sin^(10) θ + 3 sin^(8) θ+ sin^(6) θ + 2 sin^(4) θ + 2 sin^(2) θ – 2 the option are (A) 0 (B) 1 (C) 2 (D) 3

OpenStudy (anonymous):

What specific topic is this question in?

OpenStudy (anonymous):

it is the topic of polynmials

OpenStudy (anonymous):

at the end i don't get the answer

OpenStudy (anonymous):

\[\cos^6{\theta}+3\cos^5{\theta}+3\cos^4{\theta}+\cos^3{\theta}+2\cos^2{\theta}\]

OpenStudy (anonymous):

you have sum up the equation

OpenStudy (anonymous):

i m really stuck in this question